? cm^{3} cm^{3} ?? ?? selector_{m}echanism.eps, width =

140mm[GearSelectorMechanism.]Atlascargearselectormechanism.

?? P LC_{A}tlasCar.eps, width =

50mm[SiemensSIMAT ICS7− 1200P LC]SiemensSIMAT ICS7 − 1200, usedastheAtlasCarEngineControlUnit(ECU)[?]. Arduino Arduino Arduino Arduino Arduino Arduino Arduino Arduino Arduino Arduino Arduino ?? Arduino ? ?? Arduino Arduino Arduino

A B

Arduino Arduino T OP.eps[][width = 3cm]Images/V NH_{B}OT T OM.eps[V NH3SP 30− EP ackage]V NH3SP 30 − Etop(a)andbottom(b)view.NotethethreeheatslugsonthebottomviewoftheIntegratedCircuit, usedtoconductthehighcurrentsallowedbythisIntegratedCircuit. Arduino Arduino ? ?? Enable A0 A1 Q3 Q2 Q1 Q0 ??

INPUTSOUTPUTS
\overline{E}A1 A0 Q3Q2Q1Q0

Arduino Arduino ? ?? 450^{∘} pot.eps, width = 60mm[V ishaySpectrolMutiturnW irewoundP otentiometer]V ishaySpectrolMutiturnW irewoundP otentiometer[?]

Ar- duino 1080^{∘} 450^{∘} ??

V oltageInterval = 450^{∘}1080^{∘}×5 = 2.083(3) [V] (1)

??

Sensibility = 2.083(3)450^{∘} ≈ 4.62 [mV /∘] (2)

Arduino Arduino ?? pad.eps, width = 100mm[RenaulMganeadaptedaudiopad.]RenaulMganeadaptedaudiopad. Arduino ? Arduino ? a b c ?? ??

Display DigitMeaning

?? Segment_{D}isplay.eps, width =

95mm[7− SegmentDisplayF inalBoard.]7 − SegmentDisplayfinalboard. Arduino ?? Arduino ?? Arduino uno_{p}ins.eps, width = 120mm[ArduinoUNO]ArduinoUNOandtheavailablepins′functionsusedforthisproject. Arduino Arduino μ Arduino Arduino 3D_{M}ODEL.eps, width = 100mm[P artialP CB3Dmodel]P artialP CB3Dmodel, createdusingEAGLE3D.